# Demonstrating CavePI: Autonomous Exploration of Underwater Caves by Semantic Guidance

### v1

@@ 4.2 Autonomous Control
$\begin{bmatrix}u-u_{i}\\ v-v_{i}\end{bmatrix}\times\begin{bmatrix}u_{c}-u_{i}\\ v_{c}-v_{i}\end{bmatrix}=0$
@@
Ideally, the forward x-axis of the image frame aligns with the robot’s heading direction, enabling the computed heading angle to function as the primary error signal for a proportional-integral-derivative (PID) controller on the Raspberry Pi-5. The PID controller is calibrated to minimize this error, maintaining CavePI’s heading toward the designated waypoint. The controller-generated control signals are transmitted to the electronic speed controllers (ESCs) via the MAVLink communication protocol. These ESCs regulate the speed and rotational direction of the thrusters, enabling precise maneuvering and robust navigation of CavePI.
@@ 5 System Evaluation
5.1 Mechanical Stability under Hydrostatic Pressure
@@ 5.2 Caveline Following Performance
Setup. The line following experiments are initially conducted in a $2$ m×$\times 3$ m laboratory water tank with a maximum depth of $1.5$ m. As a first setup, a line is laid on the tank floor in a rectangular loop, without any depth variation; see Fig. 8. Subsequently, we add small obstacles and vertical slopes in the tank to simulate the rough underwater terrains; see Fig. 13 (a). Upon detecting the line via the down-facing camera, CavePI autonomously follows the loop while maintaining a specified depth. The system’s line-following accuracy, depth-holding capabilities, and control smoothness are evaluated and fine-tuned to prepare for field deployments.
@@
To formalize the calculation, we define several reference frames. The initial pose of the robot assumes its front axis is aligned with the caveline and its downward axis perpendicular to the caveline plane. This reference frame, denoted as {N}, has its origin at the camera. A corresponding frame {S}, with the same orientation as {N} but with its origin at the sonar, is also defined. Additionally, we define the global reference frame {G}, and the CavePI body frame {C} with its origin at the down-facing camera. The transformation matrix from any frame $a$ to $b$ is defined as ${}^{b}_{a}\mathbf{T}=[{}^{b}_{a}\mathbf{R}_{3\times 3}\,|\,{}^{b}_{a}\mathbf{t}_{3\times 1}]$. The roll, pitch, and yaw angle measurements from IMU are expressed in the global frame {G} at any time instance.
@@
${}^{I}{P}=\begin{bmatrix}u_{P}&v_{P}&1\end{bmatrix}^{T},\quad{}^{C}\boldsymbol{P}=\lambda\,K^{-1}({}^{I}P);$
@@
${}^{C}_{S}\boldsymbol{R}=\left({}^{S}_{G}\boldsymbol{R}\right)^{-1}{}^{C}_{G}\boldsymbol{R}=\left({}^{N}_{G}\boldsymbol{R}\right)^{-1}{}^{C}_{G}\boldsymbol{R};$
@@
$\begin{bmatrix}{}^{S}\boldsymbol{P}\\ 1\end{bmatrix}=\begin{bmatrix}{}^{C}_{S}\boldsymbol{R}&{}^{C}_{S}\boldsymbol{t}\\ 0&1\end{bmatrix}\begin{bmatrix}{}^{C}\boldsymbol{P}\\ 1\end{bmatrix},$
${}^{S}\boldsymbol{P}=\lambda\begin{bmatrix}a&b&c\end{bmatrix}^{T},$
@@
${{}^{S}d}={}^{C}_{S}\boldsymbol{R}_{(3,3)}\cdot{}^{C}d,~{}\text{and}~{}\lambda=\frac{{}^{S}d}{c}.$
@@
${}^{N}\boldsymbol{P}=-{}^{C}_{S}\boldsymbol{t}+{}^{S}\boldsymbol{P}$
@@ 5.3 PID Controller Tuning
To ensure stable operation in turbulent water conditions, CavePI’s PID-based Pure Pursuit controller is designed and fine-tuned through extensive experimentation. Initial trials reveal challenges, including high overshoot, difficulty maintaining depth, and instability at sharp corners. These issues are iteratively calibrated to achieve robust and reliable motion control. A grid search technique is employed to optimize the proportional gain ($K_{p}$) and differential gain ($K_{d}$) for the two onboard controllers – depth and heading. For each tested parameter pair, the robot traverses a complete 6-meter loop in the testbed at a depth of $0.35$ meters. Tracking ($\delta$) and depth errors for each trial are recorded at $0.6$ Hz. Table 2 and 3 summarize the observations of the tuning parameters combinations for heading and depth controllers, respectively. We identified that a controller tuned with $K_{p}=3.4$ and $K_{d}=0.9$ minimized the tracking error, and therefore employed these values as the heading controller gains in all subsequent experiments. Similarly, $K_{p}=600$ and $K_{d}=50$ are chosen for the depth controller.
Table 2: Tracking errors from the line following experiments are compared for various choices of $K_{p}$ and $K_{d}$ values of the heading controller. Here, $\delta$: mean tracking error (cm), $\sigma$: standard deviation.
@@
5.4 ROS Prototyping and Gazebo Simulation
The digital twin of CavePI is developed using ROS and simulated within Gazebo’s underwater environment to evaluate its line-tracking performance; see Fig. 10. As a simplified representation of the physical platform, it enables extensive pre-deployment testing of mission-critical functionalities. While the actual CavePI is positively buoyant – automatically resurfacing in the absence of control signals – the digital twin is designed to be neutrally buoyant to minimize continuous depth-control inputs. Simulating the digital twin under realistic underwater conditions allows for detailed analysis of hydrostatic and hydrodynamic forces, informing component placement for stable buoyancy and achieving near-zero roll and pitch angles when stationary. A complete sensor suite is integrated to confirm optimal camera placement, particularly for the downward-facing camera essential for caveline detection. Additionally, it incorporates a PID controller to simulate CavePI’s line-tracking functionality, serving as a platform for preliminary control algorithm design and validation. Nevertheless, final controller tuning is performed on the physical system through experimental trials due to factors such as nonlinear drag forces, caveline detection dependencies, and design parameter variations between the CAD model and the actual hardware, which cannot be replicated in simulation.
@@ 6.1 Field Trials and Experimental Setups
The guided navigation capability of CavePI was demonstrated in two distinct real-world environments: (1) shallow riverine areas ($2$ m - $6$ m depth) near springs’ outlets; and (2) deep natural underwater grottos and caves ($15$ m - $30$ m depth). The experiments included $15$ open-water trials in spring areas and $10$ trials inside underwater grottos. In the open-water trials, a $20$-meter rope line was laid in both irregular loop patterns and linear configurations along the uneven riverbed. In contrast, cave trials used an actual caveline, which varied in color, texture, and thickness. Each environment posed unique challenges. Strong currents near the spring outlets caused significant drift, particularly when navigating across or against the flow. Additionally, low-light conditions inside underwater caves hindered accurate detection of the semantic markers.
@@
Field deployments were conducted under the supervision of two support divers responsible for initiating and concluding the missions. Divers carried QR-code tags to issue visual commands to CavePI via its front-facing camera. Upon receiving the start command, the robot activated depth-hold mode and began line-following while maintaining the specified depth. A surface station operator monitored each mission remotely. For cave deployments, a tether connected the robot to the surface station to enable remote emergency intervention if necessary (see Fig. 11). The trials highlighted both the strengths and potential areas for improvement in CavePI’s perception, planning, and control mechanisms.
@@
Figure 13: A few snapshots of caveline-tracking experiments with CavePI are shown. The setups include: (a) rectangular loops and slopes in a water tank; (b) irregular shapes in Spring water, and (c) low-light underwater cave scenarios (note that the tether is for safety). Images are best viewed digitally at 2×2$\times$2 × zoom; video demonstrations can be seen here: https://robopi.ece.ufl.edu/cavepi.html.
@@ 6.2 Results and Observations
Laboratory tank tests. Sec. 5.2 details the setup and evaluation procedures for assessing CavePI’s line-following performance in laboratory tests. During these tests, CavePI operated at a depth of $0.35$ meters below the water surface. The depth controller demonstrated robust performance, maintaining the desired depth with a mean error of approximately $2$ cm. At the beginning of each trial, heading control signals were disabled for $10$ seconds to allow CavePI to stabilize at the target depth after its descent. Consequently, higher tracking errors were observed during this initial period, as shown in Fig. 9. Additionally, occasional perception failures were encountered, where the tank edges were misidentified as caveline segments (see Fig. 14 a), leading to overshooting. Despite these transient deviations, CavePI successfully completed multiple loops, demonstrating the robustness of its control strategy.
Open-water tests (Spring riverine environments). These trials were evaluated using the same methodology described in Sec. 5.2, focusing on both line-tracking and depth-regulation accuracy. Fig. 12 presents a representative result from a 10-minute open-water experiment. As expected, the field results exhibit greater variation compared to the controlled laboratory experiments due to environmental disturbances. Despite these challenges, the depth controller demonstrated robust performance, maintaining a depth within ±$\pm 10$ cm without requiring additional trial-specific parameter tuning.
In contrast, the heading controller proved less resilient under strong currents, resulting in much higher tracking errors. As illustrated in Fig. 12 b, the system experienced significant lateral drift, occasionally causing CavePI to lose the caveline. This limitation became particularly pronounced when traveling perpendicular to the current (see Fig. 15 b), as there was no active lateral control to counter crossflow. Additionally, the vehicle’s slightly back-heavy design induced a pitch-up motion during upstream travel, as shown in Fig. 15 c, further compromising the stability and navigation accuracy of the CavePI.
Low-light tests (Nighttime cave environments). These trials utilized an earlier (ablation) iteration of CavePI equipped with a three-thruster configuration, in contrast to the newly proposed four-thruster design. The three-thruster setup required precise orientation of the downward thruster to achieve proper heave motion; otherwise, thrust forces could inadvertently induce roll. Additionally, the absence of a dedicated roll-control mechanism posed challenges for stabilization. In contrast, the four-thruster configuration incorporates two thrusters for heave motion, enabling dedicated roll control and significantly improving overall stability during navigation.
@@ 7.3 Smooth 6-DOF Control
Although CavePI is a 6-DOF AUV capable of maneuvering in a 3D environment, it currently offers active control over only four DOFs – surge, heave, roll, and yaw. In future iterations, we intend to reposition its thrusters to enable control over the remaining two DOFs – pitch and sway – thereby enhancing the robot’s maneuverability in complex underwater environments. Furthermore, CavePI’s autonomous control primarily utilizes a proportional-derivative (PD) controller, which performs effectively under conditions with minimal environmental disturbances. However, this straightforward control strategy becomes unstable in more complex scenarios, such as when water currents are present. In such conditions, lower proportional gains are insufficient for adjusting CavePI’s yaw to align with the cave line, while higher proportional gains cause overshooting in the robot’s trajectory during sharp turns (see Fig. 15 d), attributed to both perception latency and the limitations of the current control method. Additionally, nonlinear (primarily quadratic) drag forces significantly impact the robot’s stability and must be incorporated into the control system design. To overcome these challenges, we are developing a more robust nonlinear adaptive control system to enhance stability. Concurrently, we are optimizing the communication between the perception and control modules to reduce latency and improve overall responsiveness.

### v3

@@ 4.2 Autonomous Control
$\begin{bmatrix}u-u_{i}\\ v-v_{i}\end{bmatrix}\times\begin{bmatrix}u_{c}-u_{i}\\ v_{c}-v_{i}\end{bmatrix}=0,$ (2)
@@
Ideally, the forward x-axis of the image frame aligns with the robot’s heading direction, enabling the computed heading angle to function as the primary error signal for a PID controller on the Raspberry Pi-5. The PID controller is calibrated to minimize this error, maintaining CavePI’s heading toward the designated waypoint. The controller-generated signals are transmitted via the MAVLink communication protocol to the electronic speed controllers (ESCs), which regulate the speed and rotational direction of the thrusters. Upon failure to detect any caveline, CavePI executes a circular search pattern to re-acquire tracking. If a line is not detected, it signals flashing lights (helps if divers are nearby); subsequently, it starts navigating towards the (last seen) arrow direction (toward the cave exit) to rediscover the caveline and/or exit the cave. Additionally, CavePI maintains a user-defined target depth by utilizing a PID controller, which processes the error betweein the current and desired depths and generates appropriate control signals to achieve stable depth regulation.
@@ 5 System Evaluation
5.1 Mechanical Stability in Hydrostatic Pressure
@@ 5.2 Caveline Following Performance
Setup. The line following experiments are initially conducted in a $2$ m×$\times 3$ m laboratory water tank with a maximum depth of $1.5$ m. As a first setup, a line is laid on the tank floor in a rectangular loop, without any depth variation; see Fig. 8. Subsequently, we add and vertical slopes in the tank to simulate the rough underwater terrains; see Fig. 15 (a). Upon detecting the line via the down-facing camera, CavePI autonomously follows the loop while maintaining a specified depth. The system’s line-following accuracy, depth-holding capabilities, and control smoothness are evaluated and fine-tuned in preparation for outdoor deployments.
@@
Figure 10: Configurations of the coordinate frames are shown at: (a) initial time $t_{0}$, and (b) an arbitrary time $t$. Note that the {C} is a (rigid) body-fixed frame attached to the robot while {N} and {S} are non-rotating (fixed orientation) frames, even though their origins translate along with the robot.
@@
To formalize the calculation, we define several reference frames; see Fig. 10. The initial pose of the robot assumes its front axis is aligned with the caveline and its downward axis perpendicular to the caveline plane. This reference frame, denoted as {N}, has its origin at the camera. A corresponding frame {S}, with the same orientation as {N} but with its origin at the sonar, is also defined. Additionally, we define the global reference frame {G}, and the CavePI body frame {C} with its origin at the down-facing camera. The transformation matrix from any frame $a$ to $b$ is defined as ${}^{b}_{a}\mathbf{T}=[{}^{b}_{a}\mathbf{R}_{3\times 3}\,|\,{}^{b}_{a}\mathbf{t}_{3\times 1}]$. The roll, pitch, and yaw angle measurements from IMU are expressed in the global frame {G} at any time instance.
@@
${}^{I}{P}=\begin{bmatrix}u_{P}&v_{P}&1\end{bmatrix}^{T},\quad{}^{C}\boldsymbol{P}=\lambda\,K^{-1}({}^{I}P);$ (3)
@@
${}^{C}_{S}\boldsymbol{R}=\left({}^{S}_{G}\boldsymbol{R}\right)^{-1}{}^{C}_{G}\boldsymbol{R}=\left({}^{N}_{G}\boldsymbol{R}\right)^{-1}{}^{C}_{G}\boldsymbol{R};$ (4)
@@
$\begin{bmatrix}{}^{S}\boldsymbol{P}\\ 1\end{bmatrix}=\begin{bmatrix}{}^{C}_{S}\boldsymbol{R}&{}^{C}_{S}\boldsymbol{t}\\ 0&1\end{bmatrix}\begin{bmatrix}{}^{C}\boldsymbol{P}\\ 1\end{bmatrix},$ (5)
${}^{S}\boldsymbol{P}=\lambda\begin{bmatrix}a&b&c\end{bmatrix}^{T};$ (6)
@@
${{}^{S}d}={}^{C}_{S}\boldsymbol{R}_{(3,3)}\cdot{}^{C}d,~{}\text{and}~{}\lambda=\frac{{}^{S}d}{c}.$ (7)
@@
${}^{N}\boldsymbol{P}=-{}^{C}_{S}\boldsymbol{t}+{}^{S}\boldsymbol{P}.$ (8)
@@ 5.3 PID Controller Tuning
To ensure stable operation in turbulent water conditions, CavePI’s PID-based Pure Pursuit controller [45, 50] is designed and fine-tuned through extensive experimentation. Pure Pursuit is a widely used geometric tracking algorithm that guides a vehicle toward a target point on a path [49] at a fixed lookahead distance. By steering the vehicle to continuously align with this moving target point, the controller ensures smooth convergence to and following of the intended trajectory, which we tuned and tested empirically.
Our initial trials reveal challenges, including high overshoot, difficulty maintaining depth, and instability at sharp corners. These issues are iteratively calibrated to achieve robust and reliable motion control. A grid search technique is employed to optimize the proportional gain ($K_{p}$) and differential gain ($K_{d}$) for the two onboard controllers – depth and heading controllers. For each tested parameter pair, the robot traverses a complete 6-meter loop in the testbed at a depth of $0.35$ meters. Tracking ($\delta$) and depth errors for each trial are recorded at $0.6$ Hz. Table 3 and 4 summarize the observations of the tuning parameters combinations for heading and depth controllers, respectively. We identified that a controller tuned with $K_{p}=3.4$ and $K_{d}=0.9$ minimized the tracking error, and therefore employed these values as the heading controller gains in all subsequent experiments. Similarly, $K_{p}=600$ and $K_{d}=50$ are chosen for the depth controller.
Table 3: Tracking errors from the line following experiments are compared for various choices of $K_{p}$ and $K_{d}$ values of the heading controller. $\delta$: mean tracking error (cm), $\sigma$: standard deviation.
@@
5.4 ROS Prototyping and Gazebo Simulation
The digital twin of CavePI is developed using ROS and simulated within Gazebo’s underwater environment to evaluate its line-tracking performance; see Fig. 11. As a simplified representation of the physical platform, it enables extensive pre-deployment testing of mission-critical functionalities. The digital twin is designed to be positively buoyant, consistent with the actual CavePI, and receives depth-control inputs to maintain a certain depth. Simulating the digital twin under realistic underwater conditions allows for detailed analysis of hydrostatic and hydrodynamic forces, informing component placement for stable buoyancy and achieving near-zero roll and pitch angles when stationary. A complete sensor suite is integrated to confirm optimal camera placement, particularly for the downward-facing camera essential for caveline detection. Additionally, it incorporates a PID controller to simulate CavePI’s line-tracking functionality, serving as a platform for preliminary control algorithm design and validation. Nevertheless, final controller tuning is performed on the physical system through experimental trials due to factors such as nonlinear drag forces, caveline detection dependencies, and design parameter variations between the CAD model and the actual hardware, which cannot be replicated in simulation. We also developed a VR interface for live interactive simulations and immersive experiments. In this interface, users wear a head-mounted display (HMD) to engage in an immersive virtual mission experience; they experience an egocentric view (i.e., robot’s perspective) as well as an external global perspective in real-time while the DT model of the robot conducts a mission. Additionally, they can switch from autonomous mode to teleoperation mode and directly control the CavePI in different virtual worlds; see Fig. 12.
@@ 5.4 ROS Prototyping and Gazebo Simulation
Figure 12: Interactive demonstration setup for CavePI’s virtual missions is shown. A user wears an HMD to experience various camera perspectives while using a handheld controller.
@@ 6.1 Field Trials and Experimental Setups
The guided navigation capability of CavePI was demonstrated in two distinct real-world environments: (1) shallow riverine areas ($2$ m - $6$ m depth) near springs’ outlets; and (2) deep natural underwater grottos and caves ($15$ m - $30$ m depth). The experiments included $15$ open-water trials in spring areas and $10$ trials inside underwater grottos. In the open-water trials, a $20$-meter rope line was laid in both irregular loop patterns and linear configurations along the uneven riverbed. In contrast, cave trials used an actual caveline, which varied in color, texture, and thickness. Each environment posed unique challenges. Strong currents near the spring outlets caused significant drift, particularly when navigating across or against the flow. Additionally, low-light conditions inside underwater caves hindered the accurate detection of the semantic markers.
Field deployments were conducted under the supervision of two support divers responsible for initiating and concluding the missions. Divers carried QR-code tags to issue visual commands to CavePI via its front-facing camera. Upon receiving the start command, the robot activated depth-hold mode and began line-following while maintaining the specified depth. A surface station operator monitored each mission remotely. For cave deployments, a tether connected the robot to the surface station to enable remote emergency intervention if necessary (see Fig. 13). These field trials’ outcomes highlighted both the strengths and potential areas for improvement in CavePI’s perception, planning, and control mechanisms.
@@
Figure 15: A few snapshots from our field trials for caveline tracking and following experiments with CavePI are shown. The setups include: (a) rectangular loops and slopes in a water tank; (b) irregular shapes in Spring riverine water, and (c) low-light underwater cave scenarios (note that the tether is only for safety and easy recovery). Images are best viewed digitally at 2×2$\times$2 × zoom; full video demonstration is submitted in the supplementary files; also available online at: https://youtu.be/9BPpB1nu98E.
@@ 6.2 Results and Observations
Laboratory tank tests. Sec. 5.2 details the setup and evaluation procedures for assessing CavePI’s line-following performance in laboratory tests. During these tests, CavePI operated at a depth of $0.35$ meters below the water surface. The depth controller demonstrated robust performance, maintaining the desired depth with a mean error of approximately $2$ cm. At the beginning of each trial, heading control signals were disabled for $10$ seconds to allow CavePI to stabilize at the target depth after its descent. Consequently, higher tracking errors were observed during this initial period, as shown earlier in Fig. 9. Additionally, occasional perception failures were encountered, where the tank edges were misidentified as caveline segments (see Fig. 16 a), leading to overshooting. Despite these transient deviations, CavePI successfully completed multiple loops, demonstrating the robustness of its control strategy.
Open-water tests: Spring riverine environments. These trials were evaluated using the same methodology described in Sec. 5.2, focusing on both line-tracking and depth-regulation accuracy. Fig. 14 presents a representative result from a 10-minute open-water experiment. As expected, the field results exhibit greater variation compared to the controlled laboratory experiments due to environmental disturbances. Despite these challenges, the depth controller demonstrated robust performance, maintaining a depth within ±$\pm 10$ cm without requiring trial-specific parameter tuning.
In contrast, the heading controller proved less resilient under strong currents, resulting in much higher tracking errors. As illustrated in Fig. 14 b, the system experienced significant lateral drift, occasionally causing CavePI to lose the caveline. This limitation became particularly pronounced when traveling perpendicular to the current (see Fig. 17 b), as there was no active lateral control to counter crossflow. Additionally, the vehicle’s slightly back-heavy design induced a pitch-up motion during upstream travel, as shown in Fig. 17 c, further compromising stability and navigation accuracy.
Low-light tests: nighttime cave environments. These trials utilized an earlier (ablation) iteration of CavePI equipped with a three-thruster configuration, in contrast to the newly proposed four-thruster design. The three-thruster setup required precise orientation of the downward thruster to achieve proper heave motion; otherwise, thrust forces could inadvertently induce roll. Additionally, the absence of a dedicated roll-control mechanism posed challenges for stabilization. In contrast, the four-thruster configuration incorporates two thrusters for heave motion, enabling dedicated roll control and significantly improving overall stability during navigation.
@@ 7.3 Smooth 6-DOF Control
Although CavePI is a 6-DOF AUV capable of maneuvering in a 3D environment, it currently offers active control over four DOFs: surge, heave, roll, and yaw. In future iterations, we intend to reposition its thrusters to enable control over the remaining two DOFs – pitch and sway – thereby enhancing the robot’s maneuverability in complex underwater environments. Furthermore, CavePI’s autonomous control primarily utilizes a proportional-derivative (PD) controller, which performs effectively under conditions with minimal environmental disturbances. However, this straightforward control strategy becomes unstable in more complex scenarios, such as when water currents are present. In such conditions, lower proportional gains are insufficient for adjusting CavePI’s yaw to align with the cave line, while higher proportional gains cause overshooting in the robot’s trajectory during sharp turns (see Fig. 17 d), attributed to both perception latency and the limitations of the current control method. Additionally, nonlinear (primarily quadratic) drag forces significantly impact the robot’s stability and must be incorporated into the control system design. To overcome these challenges, we are developing a more robust nonlinear adaptive control system to enhance stability. Concurrently, we are optimizing the communication between the perception and control modules to reduce latency and improve overall responsiveness.
8 Discussion
Our investigation identifies four primary challenges impeding effective visual servoing and safe navigation of AUVs within underwater cave environments: (1) accurate scene parsing under noisy sensing conditions, with very little ambient light; (2) robust state estimation in feature-scarce environments lacking GPS or external localization aids; (3) reliable tracking and control mechanisms to maintain visibility of navigation markers during maneuvering; and (4) the inherent logistical complexity and risks involved in field deployments for in-situ experimental validation. These challenges are further exacerbated in our experiments by the compact form factor, low inertia, and energy-efficient design of the CavePI platform, which collectively reduce motion stability in the presence of unpredictable hydrodynamic disturbances, particularly near cave entrance and spring outlets. Furthermore, the performance of our semantic-guided navigation framework hinges on the accuracy of the underlying semantic segmentation model, which is challenged by variations in waterbody types, turbidity, and backscatter commonly encountered in confined and cluttered cave systems. These findings highlight the pressing need for more robust perception and control strategies tailored to the unique operational demands of underwater caves.
To this end, the broader research community stands to benefit from our open documentation of system performance under real-world conditions for both successful outcomes and failure cases. These findings emphasize the critical need to balance perception accuracy with computational efficiency to enhance mission reliability in natural underwater environments – an aspect often underrepresented in controlled laboratory settings. Another key insight from our field deployments is the influence of external disturbances on caveline-following performance. Our results indicate that effective sway control is essential for enabling compact marine robotic systems to maintain stability and trajectory in the presence of strong currents. This observation underscores the need for further research into adaptive control strategies that can dynamically respond to environmental disturbances. Moreover, our integration of visual tags as a mechanism for human-robot interaction offers a scalable and practical alternative to tethered communication for short-range underwater tasks, advancing the potential for more flexible and intuitive operations.
Building upon this foundational system, our ongoing work focuses on advancing toward fully autonomous underwater cave exploration with greater resilience and mission safety. Recent upgrades include the integration of more powerful compute (Nvidia™ Jetson AGX Orin instead of Orin Nano), an adaptive control framework to dynamically compensate for external disturbances, and multi-modal sensor fusion [57] to enhance perception robustness and state estimation accuracy. Ongoing work is also integrating a more dense segmentation pipeline (using CaveSeg [1]) for detecting overhead structures and companion divers from sparse pointcloud (by SVIn2 [44]) to improve autonomy. Furthermore, the ability to operate in partially mapped or entirely unknown caves introduces opportunities for innovation in semantic SLAM, risk-aware path planning, and safety-critical control. These enhancements will be particularly valuable in human-robot collaborative missions, where the robot must support shared situational awareness and dynamic decision-making in real-time.